OpenStudy (anonymous):

The perimeter of a rectangle is 62. The area of the same rectangle is 234. Which of the following equations could be used to find the dimensions of the rectangle? : w(31-w)=234, w(62-2)=234, w(31+w)=234, or w(62+w)=234

OpenStudy (dbzfan836):

A:w(31-w)=234 B:w(62-2)=234 C:w(31+w)=234 D:w(62+w)=234

OpenStudy (dbzfan836):

Perimeter / 2 = 31, or one length plus one width.

OpenStudy (paki):

we have the perimeter equation: 2l+2w = 62 and solve it for l and put that value in the area equation: lw = 234 in place of l and see what we get...

OpenStudy (anonymous):

l = 31-w ?

OpenStudy (paki):

yeah A...

OpenStudy (anonymous):

Okay, I see it now. Thanks!

OpenStudy (paki):

pleasure